OpenStudy (anonymous):

A combined circuit has two resistors in parallel (10.0 ohms and 14.0 ohms) and another in series (5.0 ohms). If the power source is 9.0 volts, what is the current?

OpenStudy (kropot72):

Let the value of the two resistors in parallel be R ohms. Then\[\frac{1}{R}=\frac{1}{10}+\frac{1}{14}\ .....(1)\] Now you need to solve equation (1) to find R. Then the total combined resistance in the circuit is \[R _{total}=R+5\] the current in the circuit is given by: \[I=\frac{V}{R _{total}}=\frac{9}{R +5}\]